OpenStudy (anonymous):

what's the square root - (Square sign) 25/81 @Algebra

OpenStudy (asnaseer):

I assume you mean:\[\sqrt{25/81}=\sqrt{5 ^{2}/9 ^{2}}=\pm5/9\]

OpenStudy (anonymous):

yes